OpenStudy (anonymous):

Match each set of ordered pairs with their rise over run. Match Term Definition (4, -2), (3, 1) A) -3/7 (4, -2), (-3, 1) B) -1 (-4, -2), (3, 1) C) -3 (-4, 2), (3, 1) D) -1/7

OpenStudy (matt101):

Let's look at the first two points given. The rise over run (or the slope) is calculated by finding the change in y values (the rise) and dividing that by the change in x values (the run). So for the first two points, the slope would be (-2-1)/(4-3) = -3. Try that for the other points and you'll get your answer!

OpenStudy (anonymous):

I dont get it

OpenStudy (anonymous):

im sorry im honestly exhausted its 3 am

OpenStudy (anonymous):

@matt101

OpenStudy (matt101):

If your two points are (x1, y1) and (x2, y2) then: \[ \frac{ rise }{ run }=\frac{ y_1-y_2 }{ x_1-x_2 }\]